Write the following in logarithmic form: 2e^x=1
Answer by Boreal(15235)

ln(2)+ln(e^x)=ln(1)
ln(2)+x=0
x=-ln(2)
-----------------
alternative
e^x=(1/2)
ln of both sides
x=0-ln(2)=-ln(2)